OpenStudy (anonymous):

WILL GIVE MEDAL!!!!!!!! Simplify (3x2 - 3 + 9x3) - (4x3 - 2x2 + 16). x3 - 5x2 + 25 -x3 + x2 - 25 5x3 + x2 + 13 5x3 + 5x2 - 19

OpenStudy (anonymous):

@hugsnotughs @hugsnkisses @hellokittie4life @undeadknight26 @iGreen @iceicebaby @dan815 @Destinymasha @DangerousJesse @dillon754 @danvids @ObeyBaby_nemo @Omg_Bmth @PaulaLovesSchool13 @Preetha @perl @protected

OpenStudy (anonymous):

@AngelWilliams16 @aryandecoolest @AydenCarpenter14 @Ankh @AquaStone316 @Algorithmic @grray7 @JFraser @jonjenkins7653 @just_one_last_goodbye @justsmile531 @JReed231 @joquez14 @yayanieves @yabishshauna @Tazmaniadevil @TexasBoy123 @TylerSoSexy @TH3UPR04R187 @timden01

OpenStudy (dangerousjesse):

Combine powers and like terms.

OpenStudy (anonymous):

i'm sorry but i just need it answered please i got a few question i dont get and i need them done fast

OpenStudy (dangerousjesse):

Factor \(2\) out of \(4 x^3-2 x^2+16:\) \(3 x^2-3+9 x^3-2 (2 x^3-x^2+8)\) \(-2 (2 x^3-x^2+8) = -4 x^3+2 x^2-16:\) \(-4 x^3+2 x^2-16+9 x^3+3 x^2-3\) Grouping like terms, \(9 x^3-4 x^3+3 x^2+2 x^2-16-3 = (9 x^3-4 x^3)+(3 x^2+2 x^2)+(-3-16):\) \((9 x^3-4 x^3)+(3 x^2+2 x^2)+(-3-16)\) \(9 x^3-4 x^3 = 5 x^3:\) \(5 x^3+(3 x^2+2 x^2)+(-3-16)\) \(3 x^2+2 x^2 = 5 x^2:\) \(5 x^3+5 x^2+(-3-16)\) \(-3-16 = -19:\)

OpenStudy (dangerousjesse):

Which leaves you with...?

OpenStudy (anonymous):

dude, i'm sorry but i got no clue

OpenStudy (dangerousjesse):

Read the explanation thoroughly, then tell me the answer.

OpenStudy (anonymous):

A? idk

OpenStudy (perl):

(3x2 - 3 + 9x3) - (4x3 - 2x2 + 16 ) remove parentheses (distribute negative if necessary) = 3x2 - 3 + 9x3 - 4x3 + 2x2 - 16 = 5x^3 + 5x^2 - 19

OpenStudy (perl):

(3x2 - 3 + 9x3) - (4x3 - 2x2 + 16 ) First remove parentheses (distribute negative on second parenthetic expression) = 3x2 - 3 + 9x3 - 4x3 + 2x2 - 16 Now collect like terms = 5x^3 + 5x^2 - 19